# EGGS: Edge Guided Gaussian Splatting for Radiance Fields

### v1

@@ 1.3. Variants of Gaussian Splatting
With the advantages of computational efficiency and the resulting high quality rendered images, various Gaussian splatting methods have been developed. For example, it can be applied in street modeling (Yan et al., 2023), human head modeling (Wang et al., 2024; Zhou et al., 2024) and human body modeling (Abdal et al., 2023). In (Gong et al., 2024), isotropic Gaussian function is adopted to reduce the orientation issue. And A compression method is developed to remove the unimportant Gaussian particles, reducing the file size (Lee et al., 2023). When we are writing this paper, 2D Gaussian splatting method is developed in (Huang et al., 2024), where 2D disks are attached to a surface. Similar idea is also shown in (Guédon and Lepetit, 2023). This method is more suitable for surface representation instead of volume representation. A survey on 3D Gaussian splatting can be found in (Chen and Wang, 2024).
@@ 2.2. Our Loss Function
(14) $Loss(c,im)=(1-\lambda)\phi(u,v)\|c-im\|_{1}+\lambda D_{SSIM}(c,im)\,.$
The proposed loss function has several advantages. First, the weight function $\phi(u,v)$ is very generic and other edge indicating functions can also be adopted without changing the pipeline of the proposed method. Second, the weight function is only determined by the observed input images. It is independent from the 3D radiance field. It works in the image space rather than in the 3D object space. Therefore, it can be pre-computed without estimating the unknown 3D signal.
In contrast, some previous methods such as (Cheng et al., 2024; Huang et al., 2024) also use regularization on the estimated radiance field, for example, the smoothness of the signal or the normal consistency of the surface. However, such smoothness relies on the estimated signal or surface, which has to be updated every step during the training process.
The proposed edge weight function is fixed during the training process. Since the proposed method is a Gaussian splatting method with edge guidance. Therefore, its computational performance is exactly the same as the original 3DGS. In other words, the proposed method does not increase the computation cost, but can improve the accuracy of the radiance field. We name this method as edge guided Gaussian splatting (EGGS).
@@ 3. Experiments
Figure 2. The PSNR during the training process (the lines are smoothed for better visualization). The blue line is 3DGS and the red line is EGGS. The proposed EGGS can achieve better results on the banana data.
@@
Figure 3. The detailed difference between 3DGS and EGGS on the banana data set. As shown in the top row, the edges in the EGGS are much clearer than the edges in 3DGS. The bottom row shows that EGGS puts more particles near the edges.
@@ 3.3. PSNR Improvement
Table 1. The PSNR comparison of EGGS and 3DGS on different datasets.
@@ 4. Conclusion
In this paper, we present a simple yet effective edge guidance for the Gaussian splatting method. The proposed edge guidance is only determined by the multi view input images. And thus it does not increase the computation cost during the training and rendering stage in the Gaussian splatting method.
Moreover, as shown in the experiments, this edge guidance can improve the accuracy of the radian field about $1\sim 2$ dB, leading to a much clearer rendering results, especially at edges.
The proposed edge guidance forces the Gaussian particles to be aligned with the edges in the scene. Therefore, it will help in improving the accuracy of geometry representation. Thanks to the edge guidance, the properties such as symmetry (Gong et al., 2009) and smoothness (Gong, 2023a, b) are improved.
The proposed edge guided Gaussian splatting method can achieve higher accuracy for the scene representation and rendering. It can be applied in a large range of applications where edge information is important  (Gong et al., 2012; Yu and Orchard, 2019; Gong and Sbalzarini, 2013; Yin et al., 2019a; Gong, 2015; Yu et al., 2022; Gong and Sbalzarini, 2017; Zong et al., 2021; Gong et al., 2018, 2018; Gong and Goksel, 2019; Yin et al., 2019b; Gong and Chen, 2019; Gong et al., 2019; Gong, 2019, 2022; Yin et al., 2020; Gong and Chen, 2020; Tang et al., 2023; Gong et al., 2021a, b).

### v2

@@ 1.3. Variants of Gaussian Splatting
With the advantages of computational efficiency and the resulting high quality rendered images, various Gaussian splatting methods have been developed. For example, it can be applied in street modeling (Yan et al., 2023), human head modeling (Wang et al., 2024; Zhou et al., 2024) and human body modeling (Abdal et al., 2023). In (Gong et al., 2024), isotropic Gaussian function is adopted to reduce the orientation issue. And A compression method is developed to remove the unimportant Gaussian particles, reducing the file size (Lee et al., 2023; Fan et al., 2024). When we are writing this paper, 2D Gaussian splatting method is developed in (Huang et al., 2024), where 2D disks are attached to a surface. Similar idea is also shown in (Guédon and Lepetit, 2023). This method is more suitable for surface representation instead of volume representation. A survey on 3D Gaussian splatting can be found in (Chen and Wang, 2024).
@@ 2.2. Our Loss Function
(14) $Loss(c,im)=(1-\lambda)\|\phi(u,v)(c-im)\|_{1}+\lambda D_{SSIM}(c,im)\,.$
The proposed loss function has several advantages. First, the weight function $\phi(u,v)$ is very generic and other edge indicating functions can also be adopted without changing the pipeline of the proposed method.
Second, the weight function is only determined by the observed input images. It is independent from the 3D radiance field. It works in the image space rather than in the 3D object space. Therefore, it can be pre-computed without estimating the unknown 3D signal. In contrast, some previous methods such as (Cheng et al., 2024; Huang et al., 2024) also use regularization on the estimated radiance field, for example, the smoothness of the signal or the normal consistency of the surface. Such smoothness relies on the estimated signal or surface, which has to be updated every step during the training process.
Third, the weight function does not increase the computation cost. Its computational performance is exactly the same as the original 3DGS. In other words, the proposed method does not increase the computation cost, but can improve the accuracy of the radiance field. The proposed method is a Gaussian splatting method with edge guidance. Thus, we name this method as edge guided Gaussian splatting (EGGS).
@@ 3. Experiments
Figure 2. The PSNR during the training process (the lines are smoothed for better visualization). The blue line is 3DGS and the red line is EGGS. The EGGS can achieve better PSNR.
@@
Figure 3. The detailed difference between 3DGS and EGGS on the banana data set. The top row shows that the edges in the EGGS are much clearer than the edges in 3DGS. The bottom row shows that EGGS puts more particles near the edges.
@@ 3.3. PSNR Improvement
Table 1. The PSNR comparison of EGGS and 3DGS.
@@ 4. Conclusion
In this paper, we present a simple yet effective edge guidance for the Gaussian splatting method. The proposed edge guidance is only determined by the multi view input images. And thus it does not increase the computation cost during the training and rendering stage in the Gaussian splatting method. Moreover, as shown in the experiments, this edge guidance can improve the accuracy of the radian field about $1\sim 2$ dB.
The proposed edge guidance forces the Gaussian particles to be aligned with the edges in the scene. Therefore, it will help in improving the accuracy of geometry representation (Gong et al., 2009; Gong, 2023a, b).
The proposed edge guided Gaussian splatting method can achieve higher accuracy for the scene representation and rendering. It can be applied in a large range of applications where edge information is important  (Gong et al., 2012; Yu and Orchard, 2019; Gong and Sbalzarini, 2013; Yin et al., 2019a; Gong, 2015; Yu et al., 2022; Gong and Sbalzarini, 2017; Zong et al., 2021; Gong et al., 2018, 2018; Gong and Goksel, 2019; Yin et al., 2019b; Gong and Chen, 2019; Gong et al., 2019; Gong, 2019, 2022; Yin et al., 2020; Gong and Chen, 2020; Tang et al., 2023; Gong et al., 2021a, b; Gong, 2024).